# 0000-0002-7957-6189

### v1

@@ 1. Introduction
Emergent bilingual (EB) students in the US are acquiring English despite already communicating in a primary language other than English (García, 2009). Although EB students are the fastest-growing population in K-12 public schools, comprising 11% of the national student population, their representation in foundational computer science courses in public schools remains disproportionately low, with only 7% currently enrolled (Barr and Stephenson, 2011; Code.org et al., 2024). This gap highlights ongoing barriers to access and participation for EB students in computing classes.
For EB students, programming classes introduce a dual challenge: learning both the spoken English language and the technical English of PLs to use them in classrooms, imposing a significant cognitive load (Soosai Raj et al., 2018). Learners must process programming concepts while simultaneously navigating linguistic complexities. In the past decade, programming instruction has expanded in K-12 settings, with some states making computer science a high school (14-18 years old) graduation requirement. As programming becomes more widely available, teachers report challenges in teaching programming, specifically in classrooms with EB students. For instance, computing teachers cite struggles with their self-efficacy in programming, affecting their confidence in adopting and integrating new pedagogy and programming tools (Zhou et al., 2020; Llosa et al., 2016). Teachers with low self-efficacy may struggle to provide the appropriate scaffolding and resources to benefit EB students and allow them to participate effectively. Therefore, understanding the considerations teachers make to meet programming standards while supporting EB students is crucial.
@@
Contributing to the greater bilingual computing education research, we explore how our teacher-participants make decisions about and, thus, prioritize which PL tools are selected for general use in their programming classrooms. The research question to examine the balancing trade-offs teachers consider in this paper is: What trade-off factors shape K-12 computing teachers’ decisions when selecting and implementing programming tools to support emergent bilingual students? We identify four (4) thematic findings in response to the research question: to understand the instructional goals and the tensions between them. Broadly, the trade-offs aligned with workforce preparation without overwhelming learners, balancing cost and engagement for exploration, minimizing extraneous load through accessible representations, and scaffolding learning with familiar and intuitive syntax.
@@ 2. Prior and Related Work
The literature on bilingual computing education continues to expand within and beyond the United States. In the US, much of the research has focused on student outcomes, with a dominant emphasis on elementary and middle school students (Vogel et al., 2019; Jacob et al., 2022; Jacob and Warschauer, 2024). Internationally, studies in bilingual computing education have also centered on student outcomes, but primarily at the undergraduate level (Pal, 2016, 2016). This focus leaves a gap in research on high school students and teachers, limiting our understanding of how bilingual computing education is implemented at this critical stage.
@@ 2.1. Teacher Perceptions and Challenges in EB-Inclusive Programming Classrooms
Prior research also explains that K-12 CS teachers frequently describe their computing self-efficacy as challenging to adopt and integrating new tools into their classroom Zhou et al. (2020); Llosa et al. (2016). Teachers with low self-efficacy may struggle to offer the scaffolding and resources necessary to support EB students in succeeding in programming. On the contrary, teachers with higher computing and pedagogical self-efficacy are better equipped to design engaging activities, provide targeted feedback, and create inclusive learning environments. So, it is important to understand the considerations teachers must, or would be more willing to make, the changes that could better support EB students.
@@ 2.2. Teaching Computing in a Non-English Format
The language of instruction influences EB students’ progress, motivation, and engagement in programming education. In Botswana, Tshukudu et al. (2024) conducted a national programming outreach initiative where secondary school students were divided into two groups: one programming in English and another using a bilingual approach (Setswana and English). Their findings revealed that students in the bilingual group showed statistically significant improvements in comfort levels, suggesting that using a familiar language can create a psychologically safer and more supportive learning environment for EB students.
In another study, Raj et al. (2019) investigated the impact of incorporating Tamil in an undergraduate data structures course that used C++ as the programming language. While no statistically significant differences in student learning outcomes were observed between the English-only and bilingual (Tamil and English) groups, Tamil-speaking students in the bilingual group demonstrated higher engagement and comfort levels in classroom interactions and discussions about programming concepts (e.g., asked more questions and asked questions of higher complexity). Building on this, Molina et al. (2023) replicated that study to examine what influences language has among Spanish-speaking students in the US. They found high school students from the bilingual group (Spanish and English) were more engaged and comfortable asking questions. However, no statistically significant differences in student learning were seen among the English-only and bilingual groups. However, this further suggests that while bilingual instruction may not directly influence measurable learning outcomes, it can improve classroom dynamics and student participation.
@@ 2.3. Non-English Programming Languages
Block-based programming languages such as Scratch, Snap, Code.org’s Blockly, and MIT AppInventor offer visual approaches that reduce syntax complexity and provide immediate feedback when users drag and drop blocks to construct their programs Liao (2023); Bau et al. (2017). Some BBPLs have been localized to support other non-English languages (e.g., Scratch and Snap!), which enables EB students to code in their native language or seamlessly translanguage between English and their native tongue Vogel et al. (2019); Sarasa-Cabezuelo (2019) by toggling the ‘language’ button.
@@
Localized programming languages Bassil (2019) have also been designed to support non-English learners who want to program. These PLs are not nearly as widespread or represented in bilingual computing education literature, but they exist. Some localized PLs that have been translated include Python to Chinese Qiang et al. (2020). At the same time, French has a localized version of BASIC called LSE (Langage Symbolique d’Enseignement) Hebenstreit (1980), and Pascal, an older but well-known PL, has a Portuguese version called VisuAlg de Souza (2009). Non-English programming languages, which are foundationally designed to be in a language other than English, such as Easy-PL (Chinese), Linotte (French), Dolittle (Japanese), and PSeInt (Spanish) have served regions with limited access to English-based computing resources, fostering more equitable opportunities for students in countries across Latin America (e.g., Argentina, Bolivia, Chile, Colombia, Ecuador, Paraguay, Peru, Uruguay and Venezuela) Castrillón et al. (2022).
@@ 3.1. Recruitment
We recruited participants on Twitter, the first author’s website, K-12 computing education mailing lists, and through outreach to local school districts with previous study team connections. This is a form of purposive sampling (i.e., sampling that selects participants based on specific criteria that align with a study’s objectives) (Tongco, 2007). There were two main eligibility criteria for teachers to be considered as participants. The first was that the teacher participants must (1) be teaching a computing or programming class and (2) have EB students in their programming classrooms. We initially had 27 teacher participants who expressed interest in the study. However, a significant number (19) were excluded because they could not provide a valid school email address [^1] The decision was made following incidents where some original participants joined Zoom sessions with unprofessional or suspicious postures, which caused discomfort among the interviewers. Further investigation revealed that in some cases, the email addresses provided were personal rather than professional or were associated with outdated school systems. While some potential participants seemed genuinely interested in the research, a verifiable school email address was necessary to confirm their employment as K-12 computing teachers within the school systems under study and ensure reliable communication throughout the research process.
@@ 3.2. Design Probes and Data Collection
  • Text-based Programming Languages (TBPLs), as these require specific syntax, symbols, and keywords. These languages offer greater flexibility but a steeper learning curve, particularly for novice programmers. Participants were presented with C++ and Python as examples.
  • Text-specific Programming Languages (TSPLs) because they are designed for domain-specific tasks and have tailored interfaces. For example, Pixel Equations allows users to create custom image filters using equations to describe sections of an image, similar to graphing Dodoo et al. (2022). Colors can be defined through equations, and conditional statements allow users to test and select pixels based on their values. Participants reviewed examples from Pixel Equations (shown in both English and Spanish, Fig. 7, 7) and Charla-bots (Fig. 7).
  • Gradual programming languages (GPLs) as these introduce programming concepts incrementally. Hedy Hermans (2020), which was presented in both English and Spanish Fig. 7, 7), served as the example GPL.
@@ 3.3. Data Analysis and Validation
The interview transcripts were analyzed using reflexive thematic analysis (Braun and Clarke, 2019). The analysis began with organizing quotes and notes from the transcripts using a digital sticky-note format on Miro333www.miro.com. An affinity diagramming approach (Kawakita, 1991) was employed to cluster related excerpts and notes into initial thematic categories. Given the first author’s (ERD) background in human-centered design, the data was approached through a lens informed by the values and traditions of that field. These broad categories were iteratively refined by the second author (TNF), merging similar groups, and then we reorganized them into more focused clusters. Supporting examples were identified within the data to substantiate the emerging themes. Multiple discussions among the research team helped review and refine these themes, ensuring they accurately reflected the data. This iterative refinement process resulted in a set of prototype themes further developed into the finalized trade-offs presented in this study.
To strengthen the validity of the findings, the first author conducted follow-up conversations via Zoom with each willing participant. This process, known as member-checking (Birt et al., 2016), allowed teacher participants to review and confirm the identified themes generated by the authors. In the conversations, the teacher-participants provided additional examples and clarified statements, further strengthening the reliability of the results.
@@ 3.4. Participants
The study included K-12 computing teachers from different regions and educational settings across the United States. The participant group was evenly split by gender and included individuals self-identifying as White (4), Latiné  [^4] Latiné is a recent term to describe Spanish-speaking individuals. This term acknowledges historically problematic epistemologies. Latiné is also easier to pronounce in Spanish since it is phonologically attuned (Villanueva Alarcón et al., 2022). (3), and biracial (1). Their experience with PLs varied: half primarily used block-based languages, while the other half had experience with block-based and text-based languages. Geographically, five (5) teachers were based in Texas (covering areas along the west border, southeast, and southwest), and three (3) were in Michigan (central and southeast regions). All participating teachers reported that most of their EB students were Spanish speakers, with two (2) teachers mentioning having students who also spoke French or Korean. The predominance of Spanish-speaking EB students aligns with Spanish being the most widely spoken non-English language in the US (Dietrich and Hernandez, 2022). Additional details about participants’ teaching experience, language skills, certifications, and grade levels are provided in Table 1.
Teacher Pseudo. Total Years Teaching Race/Ethnicity Grade Level Familiarity in Teaching Programming Confidence Language(s) Spoken
@@ 4. Thematic Findings
Our teacher participants’ insights into their experiences and preferences for PL characteristics and features. The findings are subjectively what the authors understand to be important trade-offs, or priorities, that teacher participants have to decide when selecting appropriate PL tools and environments to support EB students. Note: The supporting quotes were edited to remove filler and speculation, providing clear teacher explanations.
@@ 4.1. Balancing cost and engagement for exploration
This additional time allows students to process the content at their own pace, helping them overcome language barriers and better articulate their ideas in English or their native language. This is particularly relevant when dealing with the cognitive load of translating abstract programming concepts and syntax into their native language.
@@ 6.1. Limitations
The findings of this study are drawn from a sample of eight K-12 computing teachers in the US who teach computing in classrooms, including EB students. While we interviewed teachers across various K-12 levels, geographic regions (Michigan and Texas), and course types, allowing for a range of perspectives, which likely strengthens the transferability of our findings to a broader teacher population, the sample size restricts the extent to which these findings can be generalized to the broader population of K-12 computing teachers Smith (2018). We also recognize that other teachers may have varying programming proficiency levels, experience with EB students, and familiarity with other languages their EB students may speak. Furthermore, teachers in our study had varying levels of experience with both EB students and different programming languages, which may have influenced their perspectives and the trade-offs they emphasized. A more diverse and extensive sample, encompassing a wider range of racial/ethnic backgrounds, geographic locations, and teaching contexts, would improve the generalization of these results. Future research with a larger and more diverse sample of teachers is needed to confirm and expand upon these findings and explore additional factors that may influence teachers’ decisions, such as school resources, district policies, and specific student demographics.
@@ Acknowledgments
Thank you to all the teachers for their participation, insights, and feedback on this work. We also extend our gratitude to the individuals who assisted in defining themes and engaging in meaningful discussions about interpreting the data. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-2241144. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.

### v2

@@ 1. Introduction
Emergent bilingual (EB) students in the US are acquiring English despite already communicating in a primary language other than English (Garc´ıa, 2009). Although EB students are the fastest-growing population in K-12 public schools, comprising 11% of the national student population, their representation in foundational computer science courses in public schools remains disproportionately low, with only 7% currently enrolled (Barr and Stephenson, 2011; Code.org et al., 2024). This gap highlights ongoing barriers to access and participation for EB students in computing classes.
For EB students, programming classes introduce a dual challenge: learning both the spoken English language and the technical English of PLs to use them in classrooms, imposing a significant cognitive load (Soosai Raj et al., 2018). Learners must process programming concepts while simultaneously navigating linguistic complexities. In the past decade, programming instruction has expanded in K-12 settings, with some states making computer science a high school (14-18 years old) graduation requirement. As programming becomes more widely available, teachers report challenges in teaching programming, particularly in classrooms with EB students. For instance, computing teachers cite struggles with their self-efficacy in programming, which affects their confidence in adopting and integrating new pedagogies and programming tools (Zhou et al., 2020; Llosa et al., 2016). Teachers with low self-efficacy may struggle to provide the appropriate scaffolding and resources to benefit EB students and allow them to participate effectively. Therefore, understanding the considerations teachers make to meet programming standards while supporting EB students is crucial.
@@
Contributing to the greater bilingual computing education research, we explore how our teacher-participants make decisions about and, thus, prioritize which PL tools are selected for general use in their programming classrooms. The research question to examine the balancing trade-offs teachers consider in this paper is: What trade-off factors shape K-12 computing teachers’ decisions when selecting and implementing programming tools to support emergent bilingual students? We identify four (4) thematic findings in response to the research question: to understand the instructional goals and the tensions between them. Broadly, the trade-offs align with workforce preparation, without overwhelming learners, by balancing cost and engagement for exploration, minimizing extraneous load through accessible representations, and scaffolding learning with familiar and intuitive syntax.
@@ 2. Prior and Related Work
The literature on bilingual computing education continues to expand within and beyond the United States. In the US, much of the research has focused on student outcomes, with a dominant emphasis on elementary and middle school students (Vogel et al., 2019; Jacob et al., 2022; Jacob and Warschauer, 2024). Internationally, studies in bilingual computing education have also centered on student outcomes, primarily at the undergraduate level (Pal, 2016, 2016). This focus leaves a gap in research on high school students and teachers, limiting our understanding of how bilingual computing education is implemented at this critical stage.
@@ 2.1. Teacher Perceptions and Challenges in EB-Inclusive Programming Classrooms
Prior research also explains that K-12 CS teachers frequently describe their computing self-efficacy as challenging to adopt and integrating new tools into their classroom Zhou et al. (2020); Llosa et al. (2016). Teachers with low self-efficacy may struggle to offer the scaffolding and resources necessary to support EB students in succeeding in programming. On the contrary, teachers with higher computing and pedagogical self-efficacy are better equipped to design engaging activities, provide targeted feedback, and create inclusive learning environments. So, it is important to understand the considerations that teachers must, or would be more willing to make, the changes that could better support EB students.
@@ 2.2. Teaching Computing in a Non-English Format
The language of instruction influences EB students’ progress, motivation, and engagement in programming education. In Botswana, Tshukudu et al. (2024) conducted a national programming outreach initiative, dividing students from sixteen (16) Senior Secondary Schools into two groups: one programming in English and another using a bilingual approach (Setswana and English). Their findings revealed that students in the bilingual group showed statistically significant improvements in their it programming comfort level than the English-only group, suggesting that using a familiar language can create a psychologically safer (Pedersen et al., 2022) and more supportive learning environment for the bilingual students.
In another study, Raj et al. (2019) investigated the impact of incorporating Tamil in an undergraduate data structures course that used C++ as the programming language. While no statistically significant differences in student learning outcomes were observed between the English-only and bilingual (Tamil and English) groups, Tamil-speaking students in the bilingual group demonstrated higher engagement and comfort levels in classroom interactions and discussions about programming concepts (e.g., asked more questions and asked questions of higher complexity). Building on this, Molina et al. (2023) replicated the above study to examine what influence language has among Spanish-speaking students in the US. They found that high school students from the bilingual group (Spanish and English) were more engaged and comfortable asking questions. However, no statistically significant differences in student learning were seen among the English-only and bilingual groups. However, this further suggests that while bilingual instruction may not directly influence measurable learning outcomes, it can improve classroom dynamics and student participation.
@@ 2.3. Non-English Programming Languages
Block-based programming languages such as Scratch, Snap, Code.org’s Blockly, and MIT AppInventor offer visual approaches that reduce syntax complexity and provide immediate feedback when users drag and drop blocks to construct their programs Liao (2023); Bau et al. (2017). Some BBPLs have been localized to support other non-English languages (e.g., Scratch and Snap!), which enables EB students to code in their native language or seamlessldo y translanguage between English and their native tongue Vogel et al. (2019); Sarasa-Cabezuelo (2019) by toggling the ‘language’ button.
@@
Localized programming languages Bassil (2019) have also been designed to support non-English learners who want to program. These PLs are not nearly as widespread or represented in bilingual computing education literat; however,but they exist. Some localized PLs that have been translated include Python to Chinese Qiang et al. (2020). At the same time, French has a localized version of BASIC called LSE (Langage Symbolique d’Enseignement) Hebenstreit (1980), and Pascal, an older but well-known PL, has a Portuguese version called VisuAlg de Souza (2009). Non-English programming languages, which are foundationally designed to be in a language other than English, such as Easy-PL (Chinese), Linotte (French), Dolittle (Japanese), and PSeInt (Spanish) have served regions with limited access to English-based computing resources, fostering more equitable opportunities for students in countries across Latin America (e.g., Argentina, Bolivia, Chile, Colombia, Ecuador, Paraguay, Peru, Uruguay and Venezuela) Castrillón et al. (2022).
@@ 3.1. Recruitment
We recruited participants through Twitter, the first author’s website, K-12 computing education mailing lists, and by outreach to local school districts with connections to the previous study team. This is a form of purposive sampling (i.e., sampling that selects participants based on specific criteria that align with a study’s objectives) (Tongco, 2007). There were two main eligibility criteria for teachers to be considered as participants. The first was that the teacher participants must (1) be teaching a computing or programming class and (2) have EB students in their programming classrooms. We initially had 27 teacher participants who expressed interest in the study. However, a significant number (19) were excluded because they could not provide a valid school email address [^1] The decision was made following incidents where some original participants joined Zoom sessions with unprofessional or suspicious postures, which caused discomfort among the interviewers. Further investigation revealed that in some cases, the email addresses provided were personal rather than professional or were associated with outdated school systems. While some potential participants seemed genuinely interested in the research, a verifiable school email address was necessary to confirm their employment as K-12 computing teachers within the participating school systems and ensure reliable communication throughout the research process.
@@ 3.2. Design Probes and Data Collection
  • Text-based Programming Languages (TBPLs), as these require specific syntax, symbols, and keywords. These languages offer greater flexibility but a steeper learning curve, particularly for novice programmers. Participants were presented with examples of C++ and Python.
  • Text-specific Programming Languages (TSPLs) because they are designed for domain-specific tasks and have tailored interfaces. For example, Pixel Equations allows users to create custom image filters using equations to describe sections of an image, similar to graphing Dodoo et al. (2022). Colors can be defined through equations, and conditional statements allow users to test and select pixels based on their values. Participants reviewed examples from Pixel Equations (shown in both English and Spanish Fig. 7 and Fig. 7) and Charla-bots (Fig. 7).
  • Gradual programming languages (GPLs) as these introduce programming concepts incrementally. Hedy Hermans (2020), which was presented in both English and Spanish Fig. 7 and Fig. 7), served as the example GPL.
@@ 3.3. Data Analysis and Validation
The interview transcripts were analyzed using reflexive thematic analysis (Braun and Clarke, 2019). The analysis began with organizing quotes and notes from the transcripts using a digital sticky-note format on Miro333www.miro.com. An affinity diagramming approach (Kawakita, 1991) was employed to cluster related excerpts and notes into initial thematic categories. Given the first author’s (ERD) background in human-centered design, the data was approached through a lens informed by the values and traditions of that field. These broad categories were iteratively refined by the second author, merging similar groups, and then we reorganized them into more focused clusters. Supporting examples were identified within the data to substantiate the emerging themes. Multiple discussions among the research team helped review and refine these themes, ensuring they accurately reflected the data. This iterative refinement process resulted in a set of prototype themes further developed into the finalized trade-offs presented in this study.
To strengthen the validity of the findings, the first author conducted follow-up conversations via Zoom with each willing participant. This process, known as member-checking (Birt et al., 2016), allowed teacher participants to review and confirm the identified themes identified by the authors. In the conversations, the teacher-participants provided additional examples and clarified statements, further strengthening the reliability of the results.
@@ 3.4. Participants
The study included K-12 computing teachers from different regions and educational settings across the United States. The participant group was evenly split by gender and included individuals who self-identified as White (4), Latiné  [^4] Latiné is a recent term used to describe Spanish-speaking individuals. (4), and Black (4). This term acknowledges historically problematic epistemologies. Latiné is also easier to pronounce in Spanish since it is phonologically attuned (Villanueva Alarcón et al., 2022). (3), and biracial (1). Their experience with PLs varied: half primarily used block-based languages, while the other half had experience with block-based and text-based languages. Geographically, five (5) teachers were based in Texas (covering areas along the west border, southeast, and southwest), and three (3) were in Michigan (central and southeast regions). All participating teachers reported that most of their EB students were Spanish speakers, with two (2) teachers mentioning having students who also spoke French or Korean. The predominance of Spanish-speaking EB students aligns with Spanish being the most widely spoken non-English language in the US (Dietrich and Hernandez, 2022). Additional details about participants’ teaching experience, language skills, certifications, and grade levels are provided in Table 1.
Pseudo. Total Years Teaching Race/Ethnicity Grade Level Familiarity in Teaching Programming Confidence Language(s) Spoken
@@ 4. Thematic Findings
Our teacher participants’ insights into their experiences and preferences for PL characteristics and features. The findings are based on what the authors subjectively understand to be the important trade-offs or priorities that teacher participants must consider when selecting appropriate PL tools and environments to support EB students.  [^5] Note: To improve the readability, direct quotations were edited to remove conversational filler words.
@@ 4.1. Balancing cost and engagement for exploration
This additional time allows students to process the content at their own pace, helping them overcome language barriers and articulate their ideas more effectively in English or their native language. This is particularly relevant when dealing with the cognitive load of translating abstract programming concepts and syntax into their native language.
@@ 6.1. Limitations
The findings of this study are drawn from a sample of eight K-12 computing teachers in the US who teach computing in classrooms, including EB students. While we interviewed teachers across various K-12 levels, geographic regions (Michigan and Texas), and course types, allowing for a range of perspectives, which likely strengthens the transferability of our findings to a broader teacher population, the sample size restricts the extent to which these findings can be generalized to the broader population of K-12 computing teachers Smith (2018). We also recognize that other teachers may have varying programming proficiency levels, experience with EB students, and familiarity with other languages their EB students may speak. Furthermore, teachers in our study had varying levels of experience with both EB students and different programming languages, which may have influenced their perspectives and the trade-offs they emphasized. A more diverse and extensive sample, encompassing a wider range of racial/ethnic backgrounds, geographic locations, and teaching contexts, would improve the generalization of these results. Future research with a larger and more diverse sample of teachers is needed to confirm and expand upon these findings, as well as to explore additional factors that may influence teachers’ decisions, such as school resources, district policies, and specific student demographics.
@@ Acknowledgments
We first thank the teachers for their participation, insights, and support on this work. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE-2241144. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.